OpenStudy (anonymous):

Hi, can someone please tell me how to integrate (x)(e^-x^2)

OpenStudy (raden):

int by u-sub let u = -x^2

OpenStudy (anonymous):

so I just tried it would i get something like this: -du/2 = dx?

OpenStudy (raden):

u = -x^2 du = -2x dx -du/2 = x dx

OpenStudy (anonymous):

oh whoops i forgot the x and once I integrate, would I get the following : (e^u)(x) and get (e^u)(x^2/2)

OpenStudy (anonymous):

is there any product rule involved since e^u is being multiplied by x?

OpenStudy (raden):

see your integral shoul becomes